Tom Cruise lines Inc. issued bonds five years ago at $1000 per bond. These bonds had a 30 year life when issued and the annual interest payment was then 15%. This return was in line with the required returns by bondholders at that point as described below:

Real rate of return: 5%
Inflation premium: 5%
Risk of premium 5%
Total return: 15%

Assume that five years later the inflation premium is only 3% and is appropriately reflected in the required return (or yield to maturity) of the bonds. The bonds have 25 years remaining until maturity.

New price of the bond:
Answer #1

Answer - New price of bond = $1146.60

Reason -

As after 5 years inflation is only 3% , total return will come down to 13%

Real rate of return = 5%

Inflation premium = 3%

Risk premium = 5%

Total return = 13%

And 25 years left in maturity,

Therefore,

New price of bond = coupon amount * present value annuity factor, 13%, 25 years + face value of bond * present value discount factor, 13%, 25th year.

= (1000 * 15%) * 7.32998497775 + 1000 * 0.04710195276

= 150 * 7.32998497775 + 1000 * 0.04710195276

= 1099.49774666 + 47.10195276

= 1146.60 ( rounded off )

If the solution helped, please give it a thumbs up. Thank you.